OpenStudy (anonymous):

What is the standard equation of the circle with radius 5 and center (-3,-4)

OpenStudy (alexandervonhumboldt2):

(x+3)^2+(y+4)^2=25

OpenStudy (alexandervonhumboldt2):

(a, b) center (x-a)^2+(y-b)^=r^2 where r is the radius

OpenStudy (anonymous):

Thank you!